### v1

# Assessment of the Earth orientation parameter accuracy from concurrent VLBI observations

Abstract
We have assessed accuracy of estimates of Earth orientation parameters (EOP) determined from several very long baseline interferometry (VLBI) observing programs that ran concurrently at different networks. We consider that the root mean square of differences in EOP estimates derived from concurrent observations is a reliable measure of accuracy. We confirmed that formal errors based on the assumption that the noise in observables is uncorrelated are close to useless. We found no evidence that advanced scheduling strategies with special considerations regarding the ability to better solve for atmospheric path in zenith direction applied for 1-hr single-baseline sessions have any measurable impact on the accuracy of EOP estimates. From this, we conclude that there is a certain limit in our ability to solve for the atmospheric path delay using microwave observations themselves and a scheduling strategy is not the factor that impairs accuracy of EOP determination. We determined that EOP errors vary with season, being smaller in winter and greater in summer. We found that the EOP errors are scaled with an increase in duration of an observing session as a broken power law with the power of -0.3 at durations longer than 2–4 hours, which we explain as manifestation of the presence of correlations in the atmospheric noise.
keywords:
[…]
1 Introduction
It was demonstrated in the 1980s that Very Long Baseline Interferometry (VLBI) outperforms traditional optical techniques for determination of irregularities in the Earth rotation by more than an order of magnitude [29, 43]. Regular VLBI observations dedicated to the determination of the Earth orientation parameters (EOPs) commenced on January 04, 1984 and are running since then without interruption. In order to provide a meaningful interpretation to a given parameter adjusted from measurements, its error should be characterized. It turned out to be a difficult task to assess the accuracy of the EOP from VLBI observations. VLBI does not directly measure Earth rotation. It measures voltage at receivers, which is the thermal noise with a tiny admixture of the noise from an observed extragalactic radio source. Lengthy records of voltage data — the dataset that we process for this study has over $3\cdot 10^{17}$ samples, undergo several stages of data analysis, including non-linear procedures. The last stage uses group delays as an input and estimates EOP using least squares. If we know precisely the covariance matrix of noise in group delays, we can easily arrive at dispersions of EOP estimates using the classical error propagation law. We do not know. As a coarse approximation, we form diagonal terms of the input covariance matrix from estimates of uncertainties of group delays setting off-diagonal terms to zero. It was known since 1980s [10, 24] that uncertainties of estimates of EOP and any other parameters derived that way are underestimated in a wide range from a factor of 1.2 to 10. These estimates of uncertainties are traditionally labeled as “formal” staying short from telling that these estimates of uncertainties are just wrong.
Because of a deficiency of formal errors derived with the use of a priori covariance matrix with off-diagonal terms set to zero, we have to seek alternative ways to assess the accuracy of our results. Since no other technique can estimate Earth’s spin angle more precisely than VLBI, we have to compare VLBI results against other results from independent VLBI experiments. Since the Earth orientation strongly varies with time, we cannot compare parameters determined from one epoch to parameters determined from another epoch. Thus, we arrived at a concept of concurrent independent observations, whose analysis has the potential to provide a metric of EOP accuracy more robust than formal uncertainties. In this work, we will provide results of our re-analysis of group delays of three dedicated VLBI observing campaigns in 1997.0–2000.5, in 2019–2024.5, in 2022.3–2024.7, and several other campaigns that overlapped with them in time. The goal of our study is to evaluate the accuracy of EOP determined from these concurrent observations and investigate which factors affect it.
2 Formalism of the Earth orientation parameters
Since space is three-dimensional, a rotation of a rigid body is described by three parameters. The classical formalism of rotational kinematic was established by Leonhard Euler, and these rotation angles bear his name. Considering that we estimate a small perturbation of Euler angles with respect to an a priori model, we can linearize equations of motion and write the coordinate transformation of a vector $\vec{r}_{{}_{T}}$ from the crust-fixed terrestrial coordinate system to the vector in the inertial celestial coordinate system $\vec{r}_{{}_{C}}$ in a form
$\vec{r}_{{}_{C}}=\widehat{\mathstrut\cal M}_{a}(t)\,\vec{r}_{{}_{T}}\>+\>\vec{q}_{e}(t)\times\vec{r}_{{}_{T}},$ (1)
where ^$\widehat{\mathstrut\cal M}_{a}(t)$ is the a priori matrix of the Earth rotation and $\vec{q}_{e}(t)$ is the vector of small residual Euler angles with components that we will label as $E_{1}$, $E_{2}$, $E_{3}$. Historically, polar motion and Universal Time were determined with two totally different techniques. Polar motion was determined from observations of zenith telescopes with a micrometer that was graduated with arcseconds. Therefore, a natural unit of polar motion was arcsecond. Universal Time was determined by measuring the time of passing a star through the meridian with respect to a local clock using the eye-hear method or, later, using photoelectric registration. Therefore, a natural unit of Universal Time was second. Neither method has been used for decades. Internally, partial derivatives of VLBI time delay with respect to components of the Euler rotation vector are written in radians. Therefore, natural units of EOP determined with VLBI are radians. Since we will compare $E_{1}$, $E_{2}$, and $E_{3}$, we found it is extremely inconvenient to use different units for components of the same vector and apply artificial scaling factors. Since we have to change unit scaling factors anyway, we set all of them to one and adhere SI unit for angles in our work, namely, radian. For a reference: 1 nrad = 0.206 mas = 13.8 $\mu$sec. Euler angles are related to vintage polar motion $X_{p}$, $Y_{p}$, and UT1 parameters as
$\begin{array}[]{lcl}E_{1}&=&\mu\,Y_{p}\vspace{0.5ex}\\ E_{2}&=&\mu\,X_{p}\vspace{0.5ex}\\ E_{3}&=&\kappa\,{\rm UT1},\\ \end{array}$ (5)
where $X_{p}$ and $Y_{p}$ are expressed in arcseconds, UT1 in seconds, and scaling factors $\mu=\pi/(180\cdot 60\cdot 60)\approx 4.848\cdot 10^{-6}$, $\kappa=-\pi/(12\cdot 60\cdot 60)\cdot(\Omega_{n}+\Omega_{p})/\Omega_{n}\approx-7.292115\cdot 10^{-5}$. Here $\Omega_{n}$ is the nominal Earth rotation angular velocity and $\Omega_{p}$ is the precession rate in right ascension.
Although the fact that space is three-dimensional seems firmly established, it is not uncommon to encounter a statement that there are five parameters that describe Earth rotation, for instance, in https://hpiers.obspm.fr/eop-pc/earthor/EOP.html. It was Herring et al. [11] who suggested estimating daily nutation offsets in the mid-1980s as a temporary measure because by that time the time series of VLBI experiments was too short to separate the principal nutation term with a period of 18.6 years, annual term, and precession. Nutation can be considered as a component of quasi-diurnal harmonic oscillations in the polar motion that occupies a certain part of the spectrum. Nowadays, the time span of high-quality VLBI observations is sufficient for the separation of nutation constituents, and we are in a position to return to a classical approach for estimation of the nutation term amplitudes in the frequency domain set by James Bradley in 1748 [see 28] as it has been demonstrated by Petrov [18] and now routinely used for common VLBI analysis [22]. The high-frequency spectrum of the Earth rotation exhibits a minor variability due to the free core nutation, and this variability causes estimates of the EOP spectrum to evolve with an increase in the time interval used for estimation. However, since the free-core nutation has a period of around 433 days, this variability is slow, and it is sufficient to re-evaluate the spectrum of high-frequency EOP every 2–5 months to capture this signal using a global dataset, including newly acquired data. The anharmonic constituent of EOP is much more variable and has to be estimated over a much shorter period of time, 0.5–5 days, using datasets that cover these intervals.
In the context of our work, we consider harmonic variations in the Earth’s rotation as constant for a given dataset, but Euler angles as freely varying with time. We restrict the scope of our analysis to three Euler angles estimated over intervals of 0.5 to 24 hours long.
3 Observations
We used all 24-hr geodetic multi-baseline VLBI data since 1980 April 11 through 2025 March 03 in our analysis, in total 7818 experiments. In addition, we processed data from a number of 1-hr single baseline geodetic campaigns [^1] A small number of experiments of these campaigns had more than two stations.. We split these data into several subsets of target data, reference data, and remaining data labeled as background geodetic data. We process them as well, but we report in this study only results of EOP determination from target experiments.
We designated the following multi-baseline 24-hour experiments as the target data:
  • 148 24-hr experiments of regular quad-band programs VGOS-OPS, VGOS-TEST, and VGOS-RD, collectively called in this study vo, that ran in 2019.0–2024.0. In these experiments antennas of 12–13 m diameter with the slewing speed up to $12^{\circ}/s$ dubbed as VLBI Global Observing System [VGOS; 15].
    were used. Fast slewing speed greatly simplifies schedule optimization, in particular, allowing for an increased distribution of observations among azimuth and elevation angles. The use of four bands spread over the 7 GHz bandwidth allows determination of group delay with a reported precision of 3–5 ps. The network included 13 stations. It varied from 3 to 11 stations in a given experiment, with a median of 7 stations.
  • 10 quad-band experiments at stations Mg and Ws under program VGOS-24INT-S, for brevity called s22, in 2022.3–2024.7. These experiments were scheduled as 22 consecutive 1-hr experiments using two different strategies and then concatenated. Since Ws participated in daily operational experiments under another program, the total session duration was 22 hours.
  • 81 dual-band 24-hr experiments from program CORE-A, thereafter contracted to ca in 1997.1–2000.5 [38, 39, 37] at a global network of 13 stations. It varied from 4 to 6 stations in a given experiment with the median of 6 stations.
  • 235 dual-band 24-hr experiments from concurrent NEOS-RAPID, IVS-R1, IVS-R4 programs [40] have been running since 1993.4 till present and are concurrent with one of the above-mentioned programs. For brevity, we contract program names to na and r1r4. Two 24-hr experiments are considered concurrent if the middle epochs of these experiments are within 1 hour. The network included 32 stations. It varied from 4 to 15 stations in a given experiment, with the median of 8 stations.
We also considered the following single-baseline 1-hr experiments as additional targets:
  • 3275 dual-band 1-hr experiments at stations Kk and Wz from IVS-INT-1 and IVS-INT-00 program in 1999.9–2025.2. Observations at this baseline run 3–4 days a week. We call this program KkWz.
  • 423 quad-band 1-hr experiments at stations K2 and Ws from VGOS-INT-A program in 2020.2–2025.2. Observations at this baseline run 3–4 days a week. We call this program K2Ws.
  • 133 dual-band 1-hr experiments at stations Ts and Wz from IVS-INT-2 program in 2002.8–2017.0. Observations at this baseline run 2–3 days a week till the antenna was decommissioned. We call this program TsWz.
  • 115 dual-band 1-hr experiments at stations Mk and Wz from IVS-INT-2 program in 2019.1–2025.2. Observations at this baseline run 2–3 days a week. We call this program MkWz.
All these experiments ran concurrently with 24-hr experiments from r1r4 program. Positions of stations from 1-hr programs are presented in Table 1 and their location is shown in Figure 1.
Table 1: Positions of stations participated in 1-hr single-baseline experiments.
Name Full name $\phi_{\rm geod}$ $\lambda$ $h_{\rm ort}$
deg deg m
K2 KOKEE12M 22.1264 200.3351 1151.6
Kk KOKEE 22.1266 200.3349 1160.0
Mg MACGO12M 30.6803 255.9763 1912.7
Mk MK-VLBA 19.8014 204.5445 3763.1
Ts TSUKUB32 36.1031 140.0887 44.7
Ws WETTZ13S 49.1434 12.8783 625.7
Wz WETTZELL 49.1450 12.8775 622.3
Figure 1: Station networks for different observing programs, color-coded by the number of experiments used in our analysis. Baseline from 1-hr experiments are explicitly depicted. Close telescopes (twins) are slightly offset in the maps.
The experiments listed above are not the only concurrent geodetic observations. There were other concurrent programs, but they either covered a short interval of time, for instance, a 15-day long CONT-17 program [1], or had fewer experiments. Since differences in statistics of various programs are expected to differ at a level of tens of percents, a significant dataset needs to be analyzed in order to make an inference about whether the difference in statistics are significant or not. For instance, to establish that the standard deviations of two samples that differ at 20% are significant at a 0.05 level, the F-test requires a sample size of at least 118.
3.1 Scheduling S22 campaign
During scheduling VLBI Intensive sessions, several criteria are usually taken into account to be optimized [32]. These include: (1) maximizing the total number of observations, (2) obtaining observations at a range of elevation angles to enable accurate estimation of atmospheric path delays in zenith direction, (3) ensuring a balanced distribution of observations across multiple sources, and (4) specifically including sources at low declinations and near the cusps of the mutual sky visibility [33].
While generally speaking, observations at longer baselines are more sensitive for measuring Earth’s rotation angle $E_{3}$, they also face limitations due to a reduced common visibility of the sky [33]. The $\sim\!8400$ km baseline between stations Mg and Ws presents certain geometric advantages. At this distance, observations near the zenith at one station correspond to observations at very low elevation angles at the other station (see Figure 2). This geometry is particularly beneficial for solving for residual atmospheric path delay in zenith direction, which remains the primary source of error in VLBI measurements. To effectively separate atmospheric zenith path delay from other parameters, such as station height and clock biases, it is essential to observe at a wide range of elevation angles.
Figure 2: Mutual visibility between Ws (left) and Mg (right) stations. The elevation of the other station is color-coded. The black line marks the visible horizon [35].
The s22 sessions were specifically designed to investigate improved estimation of residual atmospheric path delay in zenith direction. Each session is divided into 1-hour blocks that alternate between two scheduling strategies. Strategy A follows the conventional scheduling approach utilized in the VieSched++ scheduling software [31] and serves as a baseline reference. By contrast, strategy B is tailored to maximize the variability in elevation angles, promoting frequent alternation between high- and low-elevation observations to improve sensitivity to zenith atmospheric path delay parameters.
To implement Strategy B, sources at high elevation for each station are identified in advance. Scans are then scheduled according to the following repeating pattern: (1) a source at high elevation for Mg (corresponds to low elevation for Ws), (2) any source, (3) a source at high elevation for Ws (corresponds to low elevation for Mg), (4) any source.
The interleaved “any source” steps (2 and 4) help to maintain a balanced distribution of observations across different sources. In general, a minimum interval of 20 minutes is enforced between consecutive observations of the same source. However, due to the limited availability of high-elevation sources at each station, this restriction is relaxed to 10 minutes for those specific sources in strategy B. Additionally, to ensure adequate coverage of low-declination regions, a low-declination source is explicitly scheduled every 15 minutes in both strategies.
4 Data analysis
The first active signal amplification of the voltage at a receiver at radio frequency bands is made by the cryogenically cooled radio telescope’s front-end system. These radio frequency voltage signals of interest are typically down-converted to lower radio frequency (RF) bands for further handling. The data acquisition terminal typically comprises a digitizer for RF bands as a backend located in the operations room, which converts the analog voltage signal into a time and amplitude discrete digital representation of the frequency band-limited analog signal. These digital converter units extract a determined number of individual intermediate frequencies (IF) from each digitized RF-frequency band and convert them into the baseband channels. The individual channels are grouped as a set of channels together, parameterized by start frequency, bandwidth, and amplitude resolution. They are digitally formatted, time-tagged, and then recorded on a data storage for further processing. For example, a quad-band setup has 32 channels with 32 MHz channel bandwidth and 2-bit amplitude resolution for one linear polarization of the received radio wave. Thus, for both linear polarizations, horizontal and vertical, 64 streams of formatted data are recorded at a rate of 8 Gigisamples/s. All signal operations, from radio frequency conversion downstream to individual digital baseband channels, are coherently driven by the reference frequency of the Hydrogen maser and inherit its frequency stability.
Streams of raw voltages formatted samples are processed with the DiFX software correlator [4, 5]. It reads raw voltage samples, unpacks them from 2-bit amplitude sampling to a floating-point representation, shifts the data streams according to a coarse path delay model to some arbitrary reference points, performs Fourier transform, computes complex auto- and cross-spectra, averages them with time, and then writes the results that are called visibilities. In addition, time series with amplitude and phase of the phase calibration tones are also written into individual files for each scan and station. Results of this correlation processing are called Level 1A data [^2] See https://github.com/nasa/sgdass/blob/main/pima/doc/swin_format.pdf for specifications..
Using visibility data, we evaluated residual phase and group delay $\tau_{p}$ and $\tau_{g}$, as well as its time derivative $\dot{\tau}_{p}$, using fringe fitting software fourfit [^3] See https://www.haystack.mit.edu/wp-content/uploads/2021/01/vgos-data-processing.pdf accessed on May 16, 2025.. It combines various algorithms to calibrate for the complex bandpass of the data acquisition system and to derive group delays and delay rates using a non-linear optimization for each observation. The estimated residual group delays and phase delay rates are sought that when they are applied to phases of the complex cross-spectra time series $c_{kj}(f,t)$, the coherent sum over all spectral channels and all time epochs of a given observation reaches the maximum:
$C(\tau_{p},\tau_{g},\dot{\tau}_{p})e^{-i2\pi\,f_{0}\tau_{p}}=$ $\displaystyle\sum_{k}\sum_{j}c_{kj}\times e^{i2\pi(f_{0}\dot{\tau}_{p}(t_{k}-t_{0})\;+\;(f_{j}-f_{0})\tau_{g})}.$ (6)
In a case of quad-band data, the fringe fitting process estimates phase, ionosphere-free group delay, and the contribution of the ionosphere in a single fit. In a case of dual-band data, we process visibilities from each band separately, and then form a linear combination of group delays with coefficients selected in such a way that the frequency-dependent contribution of path delay in the ionosphere is canceled [see, for instance, 21].
Fourfit computes the uncertainties of group delays based on the achieved signal to noise ratios (SNR) assuming individual samples of visibilities are independent and uncorrelated with time and frequency using the equations 6.45 and A12.24 presented in Thompson et al. [41]:
$\sigma_{\rm gr}=\frac{\displaystyle 1}{\displaystyle 2\pi\,\rm SNR}\;\frac{\displaystyle 1}{\displaystyle\Delta f_{\rm rms}},$ (7)
where $\Delta f_{\rm rms}$ is the root mean square of IF within the band.
The ionosphere-free group delay or a combination of group delays is the basis of further data analysis.
4.1 General geodetic data analysis
We estimated parameters in a single least squares run using all geodetic data accumulated since 1980 April 11 through 2025 March 03, in total 19.8 million ionosphere-free group delay observables. Geodetic analysis is reduced to the computation of the theoretical model, forming differences between observed and theoretical delay combinations, defining the list of estimated parameters, computation of partial derivatives of group delay over these parameters, running an iterative preliminary least squares solution with outlier elimination, update of input weights based on reported group delay uncertainties by evaluation of experiment- and baseline-dependent variances that, being added in quadrature, make the ratios of weighted post-fit residuals to their mathematical expectation close to unity, and running a final single least squares solution.
The general VLBI data analysis procedure is presented in full detail in Petrov and Kovalev [22]. To avoid unnecessary repetition, we present here only highlights. We applied 3D displacements due to ocean tidal, ocean non-tidal, atmospheric pressure, and land water storage loadings provided by the International Mass Loading Service [20] and computed from the outputs of NASA Global Modeling Office numerical weather models. We applied in data reduction a priori slant path delays computed by a direct integration of equations of wave propagation through the heterogeneous atmosphere [19] using the output of NASA numerical weather model GEOS-IT which is a successor of the older model GEOS-FPIT [26, 14, 27].
The total number of estimated parameters in the final solution exceeds 5 millions. To invert a matrix of that size, we used a three-level partitioning: global parameters that were estimated using the entire dataset, local parameters that were estimated for each observing session, and segmented parameters that were estimated for each station for time interval that is shorter than an observing session.
We estimated the following global parameters: 1) positions of all the stations at the reference epoch 2000.01.01; 2) linear velocities of all the stations; 3) antenna axis offsets of most of the stations; 4) sine and cosine components of harmonic site position variations of 69 stations with a long history of observations at diurnal, semi-diurnal, annual, and semi-annual frequencies; 5) B-spline coefficients that model the non-linear motion of 29 stations that included sudden co-seismic position changes, smooth post-seismic relaxation, non-linear change of the antenna tilt, non-linear uplift due to glaciers melting, non-linear local motions, and discontinuities due to station repair; 6) coefficients of harmonic variations of the EOP vector at 890 frequencies including nutations; and 7) positions of all the sources with at least three usable observations.
When processing 24-hr experiments, we estimated Euler angles at the specified epoch and their time derivatives as local parameters. We also estimated clock breaks and baseline-dependent clock for some experiments with clock function discontinuities or baseline-dependent biases. In addition, we estimated B-spline coefficients of the 1st degree for residual atmospheric path delay in the zenith direction, tilts of the symmetry axis of the refractivity field, also known as atmospheric gradients, and clock functions for all the stations, except the one taken as a reference, as segmented parameters. Time spans of B-splines were 20 minutes for atmospheric zenith path delay, 6 hours for tilts, and 60 minutes for clock function.
When processing 1-hr experiments, we estimated Euler angle $E_{3}$ as a local parameter, clock function and residual atmospheric path delay as segmented parameters. Time spans of B-splines were 10 minutes for atmospheric zenith path delay and 20 minutes for clock function.
We also imposed constraints on estimated parameters. An exhaustive description of constraints can be found in Petrov and Kovalev [22]. Here we stress that Earth Orientation Parameters were unconstrained, and only minimum constraints were imposed on the station positions and their derivatives. The motivation for our choice of running global solutions with estimating all parameters, including EOP, in a single least squares run, rather than fixing some parameters, is that this automatically imposed consistency of the estimated station positions and EOP. It was reported in the past that fixing nutation, polar motion, or station positions to some a priori values that are considered “good” resulted in a measurable degradation of statistics because of inconsistency that a prior models [16, 6]. In addition, by solving for station positions and source coordinates, we effectively incorporate them in a solution as quantities that have uncertainties described by the full covariance matrix, while fixing them is equivalent to treating them as known exactly.
The $\chi^{2}$ per the number of degrees of freedom statistics shows that group delay uncertainties computed by the fringe fitting algorithm, $\sigma_{a}$, never match post-fit residuals. Using weights $w=1/\sigma_{a}$ results in very significant biases of estimates of standard variations of adjusted parameters. To alleviate this problem, we used weights in the form $w=1/\sqrt{\sigma_{a}^{2}+v^{2}}$, where $v$ is an extra variance computed for each experiment and for each baseline in such a way that the ratio of the weighted sum of postfit residuals to their mathematical expectation is close to unity following the approach first described in Ryan et al. [30]. The mathematical apparatus of this technique is presented in Petrov and Kovalev [22]. This additional variance $v$ turned out close to the weighted root mean square (wrms) of postfit residuals for many experiments.
4.2 Geodetic data analysis of concurrent experiments
We ran a set of global solutions, including concurrent geodetic VLBI experiments. For each solution, the processed dataset was a concatenation of three lists: background experiments, target experiments, and reference experiments. The EOP epoch of each target experiment was determined as the mean epoch of the observations used in that experiment. For 1-hr experiments, only $E_{3}$ angles were estimated, but not rates, and therefore, the EOP epoch was implied to be the middle epoch of an observing session. EOPs and their rates at the specific epoch were estimated from 24-hr reference experiments. That specific epoch was set to the EOP epoch of a corresponding target experiment. Similarly, when a target experiment had a duration of 24 hours, EOP rate and EOP at the explicitly specified epoch were estimated.
A given target experiment can be virtually shrunk by downweighting all observations outside the specified range by a factor of 1000. We used the following technique: we processed a given experiment $N$ times using data from $N$ non-overlapping segments and downweighting all other data by a factor of 1000. Figure 3 illustrates this approach. This effectively transforms a long experiment into a sequence of shorter experiments.
Figure 3: Illustration of reprocessing a given experiment four times using non-overlapping ranges of data shown with gray color and downweighting all other data.
5 Results
Here we present our results of processing different sets of concurrent experiments from global solutions that used background, target, and reference data.
5.1 Results of s22 campaign
In program s22 twenty two consecutive 1-hr VLBI experiments at a single baseline Mg-Ws were observed in 22-hr sessions that are concurrent with multi-baseline 24-hr r1 experiments. Since there were several gaps, the total count of s22 experiments is 215. We estimated $E_{3}$ angles from s22 experiments and estimated EOP and their rates from concurrent r1 experiments on the middle epoch of s22 observing sessions.
We computed biases and rms of differences in $E_{3}$ angles from each 22-hr observing session and from all the data combined. Results are presented in Table 2. Errors of both r1r4 and s22 campaigns contribute to the rms of differences. We corrected these rms by subtracting in quadrature the contribution to these differences from 24-hr programs that was set to 0.53 nrad for all the experiments. We discuss how we obtained this contribution in subsection 5.3. We call these statistics “corrected”.
Even a cursory analysis reveals a noticeable difference between experiments ran in summer months May, June, July, August, September and winter months November, December, January, February, and April. According to the classical F-test, the ratio of standard deviations, 1.85 is statistically significant at a $10^{-9}$ confidence level for a sample of this size.
Table 2: Estimates of biases and rms of differences in $E_{3}$ angles determined from 1-hr s22 experiments with respect to concurrent multi-baseline 24-hr experiment. The rms of the differences was corrected for the contribution of errors from 24-hr experiments.
Session name Session code bias rms # exp
nrad nrad
20220411_c s22401 0.34 0.52 22
20231113_a s22415 -0.06 1.28 17
20231211_a s22416 0.03 0.30 22
20240122_a s22417 0.36 0.72 22
20240219_b s22418 0.41 0.70 22
20240415_b s22420 0.27 0.97 22
20240610_b s22422 -0.12 1.31 22
20240715_b s22423 0.11 1.58 22
20240812_b s22424 0.35 1.26 22
20240916_b s22425 -0.44 1.34 22
Winter 0.25 0.78 127
Summer 0.01 1.44 88
All 0.17 1.04 215
Figure 4 illustrates the differences in $E_{3}$ angles from two observing sessions with the smallest and largest standard deviations. The differences do not exhibit a noticeable systematic pattern. The biases in $E_{3}$ angles do not exceed 60% of a standard deviation for any observing session and they are at a level of 1/6 of the standard deviation of the whole dataset.
We computed the rms of the differences for experiments scheduled according to strategy A and strategy B separately. We got 1.04 and 1.05 nrad respectively, which is totally insignificant. The differences remained insignificant when winter only or summer only experiments were used.
Figure 4: Differences in $E_{3}$ angle estimates between 22 single-baseline experiments at MgWs baseline versus estimates from concurrent ten-baseline 24-hr experiments. Left: experiment with code s22416 on December 13, 2023. Right: experiment with code s22423 on July 15, 2024.
5.2 Results from regular 1-hr observing campaign at KkWz
The goal of 1-hr campaigns is to provide $E_{3}$ angle estimates with a short latency. For instance, the goal of IVS-INT-00 campaign is to have a latency of 36 hours, and this explains why these campaigns are called intensive. To meet these requirements, the experiment duration is reduced to one hour and the number of stations is reduced to two. Occasionally, a third and fourth station is added. A significant number of these experiments overlap with r1r4 24-hr observing sessions. We set EOP reference epochs for corresponding r1r4 experiments to the middle epoch of 1-hr experiments and derive $E_{3}$ angles at exactly the same epochs from 1-hr and 24-hr experiments. We computed statistics of the differences and applied an iterative procedure starting with the data point with the greatest $|x|/$rms ratio and consecutively eliminating such points and re-computing the rms till no data points with this ratio greater than 3 remained. The number of outliers does not exceed 2%. The statistics of the differences corrected for outliers, as well as the mean values of the corresponding formal errors are shown in Table 3.
Table 3: Differences in the $E_{3}$ angle estimates between regular single-baseline 1-hr experiments versus estimates from the concurrent multi-baseline 24-hr experiments for the same epoch. The rms of the differences were corrected for the contribution of errors from 24-hr experiments. The last column provides the arithmetic mean of formal errors from our solutions.
Baseline # exp bias rms mean $\sigma_{f}$
nrad nrad nrad
K2Ws 409 0.29 1.30 0.59
MgWs 29 -0.17 1.22 0.62
MgWs 215 0.17 1.04 0.66
TsWz 107 0.21 1.45 0.99
IsWz 65 0.74 1.43 0.86
MkWz 115 0.04 1.21 0.72
HnMk 58 0.75 1.60 1.36
KkWz 2905 -0.22 1.50 1.26
MgWs denotes estimates from s22 program, while MgWs denotes estimates from other programs.
We computed the distribution of residuals divided by scaled uncertainties and show the plot for experiments at baseline K2Ws and KkWz in Figure 5. For comparison, we also plot the Gaussian distributions with the first moments equal to the bias and with fitted second moments. We see that the distribution of normalized differences of $E_{3}$ angles determined from observations at KkWz baselines has a positive excess kurtosis, 0.84, i.e. that distribution has heavier tails than the normal distribution in the regions of normalized residuals that are by a modulo above 2.5. This is a common feature of empirical distributions. The fitted second moment, 2.60 for K2Ws and 1.18 of KkWz baseline is within 5–10% of the ratio of the rms to the mean formal uncertainty. We should note that baselines K2Ws and KkWz are geometrically almost identical. The main difference is the recorded bandwidth and therefore, the reported group delay uncertainties.
Figure 5: The normalized distributions of the $E_{3}$ angle differences derived from analysis of experiments at K2Ws (left) and KkWz baselines (right) divided by formal uncertainties. The blue smoothed lines show Gaussian distributions with the second moments 2.60 and 1.18 respectively.
We show in Figure 6 the differences in $E_{3}$ angles from the INT-INT-1 campaign that has been running for over 25 years. We do not see a significant systematic pattern in the plot of residuals. Then we computed the rms over six summer months May, June, July, August, September, October and over six winter moths. The rms of differences are 1.57 and 1.39 nrad respectively. Since the number of winter and summer concurrent experiments is over 1400, the F-test shows that this difference is significant at a $10^{-5}$ level. Winter IVS-INT-1 experiments are more accurate than summer experiments. Summer experiments on average add in quadrature 0.78 nrad variance.
Figure 6: Differences in the $E_{3}$ angle estimates derived from processing data at single-baseline 1-hr experiments at KkWz baseline versus concurrent multi-baseline 24-hr r1 experiments as a function of time.
We performed a similar summer/winter test for VGOS-INT-A campaign at baseline K2Ws. Both ends of baseline K2Ws and KKWz are located within less than 100 meters of each other. The corrected rms of the differences are 1.11 nrad for winter and 1.42 nrad for summer. According to F-test these differences are significant at a $6\cdot 10^{-4}$ level. Summer experiments on average add in quadrature 0.88 nrad variance.
5.3 Results from regular concurrent 24-hr VLBI campaigns
We have processed data from 24-hr concurrent VLBI campaigns in a similar way. We estimated the EOP and their rates on epochs that are the arithmetic mean of the middle epochs of each individual experiment. We computed the mean value of the differences (bias) and their rms. Since a priori we do not have knowledge that EOP estimates from one of the campaigns vo versus r1r4 or ca versus na are substantially different, we assume they are equally accurate and uncorrelated. Since observing stations of concurrent networks are typically separated by hundreds of kilometers and more, we consider that atmospheric path delay is expected to be uncorrelated at these distances to a great extent. Therefore, we divide the rms by $\sqrt{2}$ and report in Table 4 the scaled rms which we view as a measure of the average accuracy of concurrent vo/r1r4 and ca/na campaigns.
Table 4: Accuracy of EOP determination from comparison of 24-hr concurrent observing sessions under ca/na and vo/r1r4 programs. The rms of differences were divided by $\sqrt{2}$.
Campaign Comp bias rms units # points
ca/na $E_{1}$ 0.30 0.57 rad 82
ca/na $E_{2}$0.06 0.83 rad 82
ca/na $E_{3}$ 0.03 0.59 rad 82
ca/na $\dot{E}_{1}$ 0.59 2.26 $10^{-14}$ rad/s 82
ca/na $\dot{E}_{2}$0.30 2.23 $10^{-14}$ rad/s 81
ca/na $\dot{E}_{3}$0.43 1.45 $10^{-14}$ rad/s 82
vo/r1r4 $E_{1}$ 0.18 0.52 rad 155
vo/r1r4 $E_{2}$ 0.17 0.71 rad 156
vo/r1r4 $E_{3}$0.20 0.44 rad 157
vo/r1r4 $\dot{E}_{1}$0.02 1.45 $10^{-14}$ rad/s 155
vo/r1r4 $\dot{E}_{2}$ 0.07 1.53 $10^{-14}$ rad/s 156
vo/r1r4 $\dot{E}_{3}$0.04 1.11 $10^{-14}$ rad/s 157
We see the biases are less than $0.5\sigma$, i.e. insignificant. It is worth noting that although the rms of differences of $E_{1}$ and $E_{2}$ from vo/r1r4 campaigns is smaller than from ca/na campaigns that ran 25 years earlier, this difference is not statistically significant even at a 0.20 level. The differences in $E_{3}$ accuracies are significant at a 0.002 level and in $\dot{E}_{3}$ are significant at a 0.005 level: vo/r1r4 results are more accurate. The differences in rms of $\dot{E}_{1}$ and $\dot{E}_{2}$ are significant at $10^{-5}$.
$E_{1}$, $E_{2}$, and $\dot{E}_{3}$ are regularly estimated during analysis of GNSS data by a number of groups. This prompted us to use the three corner hat method [8] for evaluation of uncertainties of these parameters from GNSS, vo, and r1r4 campaigns. Indeed, when three time series are uncorrelated, we can derive the variances of each series from the variances of their differences. We took daily time series of polar motion and lengths of day from the IGS final series, converted them to $E_{1}$, $E_{2}$, and $\dot{E}_{3}$, interpolated to the epoch of observations, computed the rms of the differences, and then derived standard deviations of each three series. Results are presented in Table 5. We are aware that the three corner hat method is susceptible to the presence of weak correlation and systematic errors. We computed a smoothing spline at 9 equally distributed knots to fit the difference between the series and removed a small systematic pattern, but this did not change the results noticeably.
We see that the accuracy of polar motion and $\dot{E}_{3}$ from GNSS has improved by a factor of 2.5 over 25 years since 1997–2000.5. Our accuracy estimate, 0.20 nrad for $E_{1}$ and $E_{2}$, is 50% higher than the estimate of Ray et al. [25] (0.12–0.15 nrad). The EOP accuracy from VLBI also improved, but the improvement is not that strong: 28 to 68%. The results of our comparison shows that accuracy of $E_{1}$ and $E_{2}$ from r1r4 sessions is better than both accuracy from old ca/na campaigns and from vo campaign. According to the F-test, this accuracy difference is significant at a level of 0.002. The difference in uncertainties of $\dot{E}_{3}$ from these campaigns is significant at the 0.01 level. Our tests confirm that accuracy of $E_{1}$ and $E_{2}$ from GNSS was on par with VLBI in late 1990s, but it improved by a factor of 2–3 by the 2020s.
Table 5: Upper table: three corner hat analysis of rms of the estimates of $E_{1}$, $E_{2}$, $\dot{E}_{3}$ from GNSS, ca, and na VLBI programs over 1997.0–2000.5. Lower table: the same analysis for GNSS, vo, and r1r4 VLBI programs over 2019.0–2024.0. Column 2: rms of differences between GNSS results (IGS Final) versus results from VLBI ca or vo experiments. Column 3: rms of differences between GNSS results versus results from VLBI na or r1r4 experiments. Column 4: rms of differences between VLBI results from ca versus na experiments or from vo versus r1r4 experiments. Columns 5–7: rms of estimates of $E_{1}$, $E_{2}$, $\dot{E}_{3}$ derived by the three hat corner algorithm. Columns 5: rms of GNSS results. Columns 6: rms of results from VLBI ca or vo experiments. Columns 7: rms of results from VLBI na or r1r4 experiments. Columns 8: unit.
(1) (2) (3) (4) (5) (6) (7) (8)
Comp $\sigma_{\rm gnss,ca}$ $\sigma_{\rm gnss,na}$ $\sigma_{\rm ca,na}$ $\sigma_{\rm gnss}$ $\sigma_{\rm ca}$ $\sigma_{\rm na}$ unit
$E_{1}$ 0.71 0.65 0.66 0.50 0.51 0.42 $10^{-9}\>\;$ rad
$E_{2}$ 0.63 0.67 0.73 0.40 0.49 0.54 $10^{-9}\>\;$ rad
$\dot{E}_{3}$ 3.26 2.93 2.10 2.72 1.80 1.08 $10^{-14}$ rad/s
Comp $\sigma_{\rm gnss,vo}$ $\sigma_{\rm gnss,r1r4}$ $\sigma_{\rm vo,r1r4}$ $\sigma_{\rm gnss}$ $\sigma_{\rm vo}$ $\sigma_{\rm r1r4}$ unit
$E_{1}$ 0.52 0.38 0.60 0.16 0.46 0.34 $10^{-9}\>\;$ rad
$E_{2}$ 0.62 0.40 0.65 0.24 0.57 0.32 $10^{-9}\>\;$ rad
$\dot{E}_{3}$ 1.45 1.34 1.31 1.17 1.04 0.84 $10^{-14}$ rad/s
5.4 Results from 1-hr blocks of concurrent 24-hr VLBI campaigns
In subsection 5.1 we explained how we compared a dedicated campaign s22 that contained 22 1-hr schedules in one observing session. These 1-hr schedules were designed independently. Here we expand this approach to 24 hours experiments. These experiments were scheduled as single 24-hr blocks and were not designed to be processed as 1-hr separate blocks. Nevertheless, we consider that it would be instructive to analyze concurrent experiments that way.
We reprocessed two concurrent campaigns, ca/na and vo/r1r4. Results are shown in Table 6. It is worth comparing these rms of the differences against Table 4 with the results of processing whole 24-hr data chunks. As expected, shrinking the dataset by a factor of 24 caused an increase in rms of EOP estimates, in particular, a factor of 3.7 for ca/na and 2.8 for vo/r1r4 for $E_{3}$. But this reduction of rms is short of a naive $\sqrt{24}\approx 4.9$ factor.
Table 6: Accuracy of EOP determination from comparison of 24-hr observing sessions under programs vo/r1r4 and ca/na from 1-hr segments. The rms of differences were divided by $\sqrt{2}$.
Campaign duration # seg $E_{1}$ $E_{2}$ $E_{3}$
hour nrad nrad nrad
ca/na 1 1763 3.53 3.90 2.16
vo/r1r4 1 3348 1.61 1.81 1.23
We found this phenomenon deserves further investigation. We ran a battery of solutions and varied segment duration from 1 hour to 23 hours with a step of 0.5 hour in a range of 1 to 6 hours and with a step of 1 hour for segments longer than 6 hours. The dependence of the rms of differences in EOP estimates is a monotonously decaying function of the segment length. The EOP accuracy as a function of the segment length for ca/na and vo/r1r4 programs is shown in Figure 7.
Figure 7: Rms of errors in $E_{1}$ (middle green line), $E_{2}$ (upper blue line), $E_{3}$ (red bottom line) as a function of segment duration. The left plot shows results for ca/na campaigns. The right plot shows results for vo/r1r4 campaigns.
We tried to approximate this dependence with a smooth curve. Since we do not know a priori its functional dependence, adhering the Occam’s principle, we were looking for the simplest form, such as an exponential decay or a power law. Neither of these simple functions fit the data. We resorted to the broken power law that is commonly used in astrophysics in the form of
$f(t)=\left\{\begin{array}[]{l}t^{\beta_{1}}\enskip{\rm if}\enskip t>b\vspace{1ex}\\ t^{\beta_{2}}\enskip{\rm if}\enskip t\leq b\end{array}\right.,$ (10)
where $b$ is a breaking point. We found numerically $b$, $\beta_{1}$, and $\beta_{2}$ that minimize the sum of squares of residuals on a grid using a brute force approach. The results for $E_{1}$ and $E_{3}$ angles are shown in Figure 8.
Figure 8: Rms of errors in estimates of $E_{1}$ (left) and $E_{3}$ (right) angles as a function of experiment duration from vo/r1 comparison (green dots). The middle blue line shows the best fit of the broken power laws, while the bottom black line shows the power law with $\beta=-0.50$, representing the expectation based on uncorrelated observations.
We see that for segments shorter than 2–3 hours, the rms of EOP differences decreases quite fast with an increase in the amount of data in a rate that is comparable with a square root of the number of data samples. For longer segments, the rms of EOP differences decreases much slower than the square root power law. It follows from the theorem of Gauss-Markov that if the noise is uncorrelated, then the variance of least square estimates for a given set of estimated parameters is reduced as square root of the number of observations. Therefore, we interpret Figure 8 as a direct evidence of a violation of the assumption that the noise is uncorrelated.
In order to confirm this hypothesis, we ran another set of solutions and added uncorrelated, Gaussian, zero-mean noise produced by the random noise generator with specified second moments. When we add a noise with known characteristics, we dilute the properties of the intrinsic noise. We expect that, according to the Gauss-Markov theorem, the dependence of the EOP accuracy as a function of segment length should asymptotically converge to the power law -0.5 when the amount of the injected uncorrelated noise becomes dominant. Results are presented in Table 7. We see that indeed, adding uncorrelated noise made estimates of $\beta_{2}$ closer to -0.5. This pattern is seen more clearly for $E_{3}$ angle than for $E_{1}$ and $E_{2}$. Parameter $\beta_{2}$ converges faster for ca/na experiments compared to vo/r1r4.
Table 7: The estimates of the model of scaling EOP errors with the observing session duration with a broken power laws with powers $\beta_{1}$ at spans less than the breaking point $b$ and $\beta_{2}$ after that. Computation is done for two concurrent campaigns vo/r1r4 and ca/na in four modes: when original data are used and when uncorrelated white noise with the second moment 20, 50, and 100 ps was added.
Component $\beta_{1}$ $b$ $\beta_{2}$ Component $\beta_{1}$ $b$ $\beta_{2}$
$E_{1}$ vo/r1r4 -0.478 2.16 -0.276 $E_{1}$ ca/na -0.895 1.70 -0.510
$E_{2}$ vo/r1r4 -0.460 2.48 -0.303 $E_{2}$ ca/na -0.765 2.13 -0.444
$E_{3}$ vo/r1r4 -0.430 4.50 -0.292 $E_{3}$ ca/na -0.693 2.87 -0.331
$E_{1}$ vo/r1r4 + 20ps -0.570 1.74 -0.308 $E_{1}$ ca/na + 20ps -0.792 2.05 -0.517
$E_{2}$ vo/r1r4 + 20ps -0.481 2.55 -0.330 $E_{2}$ ca/na + 20ps -0.744 3.30 -0.427
$E_{3}$ vo/r1r4 + 20ps -0.526 2.20 -0.345 $E_{3}$ ca/na + 20ps -0.692 2.68 -0.392
$E_{1}$ vo/r1r4 + 50ps -0.506 3.19 -0.311 $E_{1}$ ca/na + 50ps -0.782 3.00 -0.531
$E_{2}$ vo/r1r4 + 50ps -0.502 2.61 -0.344 $E_{2}$ ca/na + 50ps -0.734 3.50 -0.534
$E_{3}$ vo/r1r4 + 50ps -0.575 2.85 -0.374 $E_{3}$ ca/na + 50ps -0.709 3.07 -0.505
$E_{1}$ vo/r1r4 + 100ps -0.563 3.81 -0.341 $E_{1}$ ca/na + 100ps -0.763 4.21 -0.539
$E_{2}$ vo/r1r4 + 100ps -0.821 1.71 -0.445 $E_{2}$ ca/na + 100ps -0.806 2.06 -0.606
$E_{3}$ vo/r1r4 + 100ps -0.668 1.87 -0.488 $E_{3}$ ca/na + 100ps -0.771 2.19 -0.628
5.5 Results from 0.5-hr blocks of 1-hr VLBI campaigns
In a similar way, as we split 24-hr observing sessions into 1-hr blocks, we split the 1-hr observing sessions at K2Ws baseline into two 0.5-hr blocks. We selected that campaign because, on average, there are 60 observations in each 1-hr experiment, and the experiments were designed to observe low and high elevation sources every 10–30 minutes. We computed $E_{3}$ angles using only the first 30 minutes of data and the last 30 minutes of data. Then we computed $E_{3}$ angles from concurrent 24-hr observations, formed the differences, and computed the rms. The rms corrected for errors in $E_{3}$ from 24-hr observing sessions over 346 epochs is 1.42 nrad. In comparison, when the full 1-hr blocks in KsWs were used, the resulting corrected $E_{3}$ rms is 1.30 nrad. An increase in the number of observations used in a solution by a factor of 2 caused a reduction of errors by 9%. However, according to the F-test, this difference is statistically insignificant even at a 10% level.
We also computed statistics of $E_{3}$ angles for the same experiments from two 0.5-hr blocks from the same experiment against each other by forming differences $E_{3}(t_{2})-E_{3}(t_{1})-\dot{E}_{3}(t1)\;(t_{2}-t_{1})$. We took $\dot{E}_{3}(t1)$ from results of processing 24-hr experiments. We found that the rms of the differences is 1.47 nrad.
6 Discussion
Here we provide interpretation of our results.
6.1 Outcomes of different scheduling strategies
Within the s22 experiments, we explicitly tested an observing strategy that we expected would provide a more precise evaluation of the atmospheric path delay from the observations themselves, because we included observations of sources at high and low elevation angles in a rapid sequence over short periods of time. We generated simulated datasets from schedules and added the noise based on the mathematical model of atmospheric turbulence and its code developed by Pany et al. [17]. We ran a battery of least square solutions of the simulated datasets with a diagonal weight matrix, generated time series of simulated $E_{3}$ estimates and computed their rms: 1.24 and 1.21 nrad for scheduling strategies A and B respectively.
Our analysis of scheduling blocks designed with an advanced strategy revealed no significant improvement compared to results from processing scheduling blocks designed according to the traditional strategy. Does it mean that estimation of the atmospheric path delay plays no role? To answer these question, we ran two additional solutions and processed 215 s22 sessions in two modes: 1) we estimated a single atmospheric path delay for 1-hr blocks and 2) we did not estimate zenith path delay. Results of these two auxiliary solutions, as well as the reference solution, are presented in Table 8.
Table 8: Impact of estimation of the atmospheric path delay in zenith direction on $E_{3}$ estimates from s22 experiments. The first row corresponds to the main solution with estimation of path delay in zenith direction with a linear spline with segment lengths of 10 minutes. The second row corresponds to the case when one parameter of atmospheric path delay in zenith per a 1-hr block was estimated. And the last row corresponds to a case when atmospheric path delay was not estimated. The third column labeled accuracy is the rms of $E_{3}$ angle estimates from concurrent VLBI experiments.
Solution bias accuracy formal error postfit rms
nrad nrad nrad ps
10 min segments 0.22 1.04 0.74 25.61
60 min segments 0.19 1.04 0.69 29.89
no estimation 3.35 1.46 0.66 56.31
We see estimation of the atmospheric path delay in zenith direction does reduce postfit statistics, eliminates biases, and improves the accuracy, but degrades formal errors. However, estimating atmospheric path delay every 10 minutes in the reference solution makes no differences with respect to estimating it every 60 seconds.
We interpret these results as a manifestation of a certain limit of our ability to estimate atmospheric path delay from observations themselves. Atmospheric turbulence does not vanish at scales less than 60 minutes, but modeling slant atmospheric path delay with the zenith path delay scaled by the atmospheric mapping function shows it does. To explain this contradiction, we suggest that modeling residual atmospheric path delay as a product of the zenith path delay and the mapping function is not adequate at scales shorter than one hour.
It is worth noting that if the minimum formal uncertainty were chosen as the criterion for selection of the optimal estimation strategy, one could select an approach of not estimating atmospheric path delay in zenith direction, which in fact, provides the highest bias and highest rms of differences between $E_{3}$ estimates from concurrent VLBI experiments.
6.2 Seasonal accuracy variability
Our analysis of 1-hr experiments at MgWz and KkWz reveals a strong seasonality in the accuracy of $E_{3}$ angles estimates: the accuracy is worse in summer than in winter by the root sum squared 1.2 and 0.7 nrad for these two baselines, and this difference is highly statistically significant. However, the overall picture is not that clear. We show in Figure 9 the seasonal pattern by plotting the difference in $E_{3}$ angle estimates from K2Ws baseline with respect to the multi-baseline 24-hr experiments as a function of a fractional part of the year defined as the ratio of the day of a year to 365. The fractional part of a year 0.0 corresponds to the 1st January. The pattern shows several shallow maxima and minima.
Figure 9: Differences in the $E_{3}$ angle estimates between single-baseline experiment at K2Ws baseline versus concurrent multi-baseline 24-hr r1 experiments as a function of a fractional part of a year.
We noticed in 1990s that the scatter of the postfit residuals from geodetic observations at very sensitive Very Long Baseline Array (VLBA) antennas is noticeably greater in summer than in winter. The left plot in Figure 10 demonstrates this pattern for modern quad-band vo VLBI experiments. The rms of postfit residuals in July is a factor of 1.6 greater than in January. For comparison, we show a plot of zenith path delay at Wz stations also as a function of the fractional part of a year.
Figure 10: Left: the rms of postfit residuals of 24-hr vo experiments as a function of a fractional part of a year. Right: zenith wet path delay at station Wz as a function of a fractional part of a year.
Putting together these two findings, we conclude that the dominant error source in EOP estimates is due to the residual contribution of the atmospheric path delay that was not captured by the estimation of zenith path delay and tilts of the refractivity field symmetry axis. Formal uncertainties of group delays in quad-band experiments are at a level of 3–5 ps all year round, while the scatter of postfit residuals varies in a range of 15–25 ps. There should be some source of these errors that is greater in summer than in winter. Path delay is greater in summer. If the atmospheric path delay is entirely captured by the estimation process, an increase in wet path delay would not affect residuals. When estimating zenith path delay and tilts captures only some share of the total signal, postfit residuals will depend on the total path delay. The unmodeled part of path delay will affect estimates of other parameters, such as EOP. The evidence we presented supports this explanation.
Seasonality in errors is the most profound when most of the network stations are located in mid-latitudes in the same hemisphere. Observations at a network that includes equatorial stations (no winter) or stations in different hemispheres (winter at one end of a baseline and summer at the other end) will exhibit less seasonality.
6.3 EOP accuracy as a function of an observing session duration
Our extensive analysis of the impact of observing session duration on EOP estimation revealed an interesting pattern from the comparison of vo and r1r4 experiments. When we increase session duration from 1 to 2–4 hours, the EOP accuracy improves with the power law exponent [-0.43, -0.46]. After reaching the breaking point, the rms of differences decreases with the power law exponent [-0.28, -0.30], much lower than the power law of -0.5 that describes the dependence of uncertainties of estimated parameters in the presence of uncorrelated noise. We treat this power law as a strong empirical evidence of the presence of correlation in noise.
What is the origin of this correlated noise? We consider the noise in group delays that contributes to right-hand sides of observation equations has three components: 1) thermal noise in the receiver; 2) instability of path delay in the VLBI signal chain hardware; 3) contribution from the unmodeled phenomena, either deterministic or stochastic. The first contribution is due to radiation from individual molecules of receivers, which number is comparable to the Avogadro number ($6\cdot 10^{23}$) and therefore, is totally uncorrelated. In a similar way, hardware errors are expected to be weakly correlated. But unmodeled phenomena are expected to be correlated. The power law exponent [-0.28, -0.30] implies that the overall contribution of all noise sources combined is highly correlated. As we saw, estimation of the atmospheric path delay in zenith direction retrieves the residual path delay only partly. The remaining path delay propagates to the right hand side. Considering that EOP errors have a strong seasonal variability, we connect these unmodeled phenomena with the contribution of residual atmospheric path delay that was only partly captured by the estimation process.
When we injected an uncorrelated noise with a different variance, the power law is by modulo increasing, approaching -0.5 with an increase of the variance of the injected noise. This confirms that the power law of the dependence of EOP accuracy on session duration is indeed due to properties of the dominant noise contribution: when we added uncorrelated noise, we diluted the intrinsic noise.
The power law of the dependence of EOP accuracy on session duration is noticeably different for EOP derived from vo/r1r4 data and derived from ca/na data. Two factors play a role. Firstly, the average reported uncertainty of group delay from vo experiments is 4 ps, while it is 20 ps for other observing sessions. That implies that the contribution of the thermal and hardware noise that is weakly correlated is minimal, and the share of the atmospheric noise is maximal. Our test with injection of an additional uncorrelated noise suggests that this factor contributes. Secondly, the average number of observations in vo 24-hr experiments is 9110, compared with 4328 observations in r1r4 experiments, 1204 observations in ca campaign, and 988 observations in na campaign. Our results indicate that the power law exponent depends not only on session duration, but on the total number of observations: the more observations, the flatter the power law.
The fact that the dependence of EOP accuracy on segment duration cannot be described by a single power law implies that there is another factor that impacts accuracy of EOP derived from processing segments that are shorter than 2–4 hours. We should note that these experiments were designed as 24 hour blocks, and we artificially split them into shorter blocks. When a segment becomes too short, separation of estimated parameters, such as EOP, clock functions and atmospheric path delays in zenith direction is getting worse. Observations at high and low elevations need to be present in a segment for a reliable estimation of atmospheric path delay. A schedule optimized for a 24-hr block guarantees that such observations are present in a 24-hr block, but it does not guarantee that such observations are present in any 1, 1.5, 2, or 2.5 hour long blocks. As a result, atmospheric path delay errors affect EOP estimates from processing shorter segments to a greater extent than they affect EOP from processing longer segments.
Our results of processing 0.5 hr segments add an additional argument. The rms of $E_{3}$ angles decreased by a factor of 1.09 ($2^{-0.13}$) when the segment duration increased by a factor of 2. At the same time, the rms of $E_{3}$ angles decreased by a factor of 1.37 ($2^{-0.45}$) when segment duration increased by a factor of 2 from 1 hour to 2 hours in vo/r1r4 observing campaigns. We attribute this discrepancy to differences in scheduling strategy: vo/r1r4 were not designed to run as short segments, but 1-hr experiments were.
VLBI data that were used for deriving angles $E_{3}(t_{1})$ and $E_{3}(t_{2})$ from 0.5 hr blocks at $t_{1}$ and $t_{2}$ epochs are independent. The thermal noise in these data is independent and uncorrelated. If the only contribution to the noise were the thermal noise in receivers, then rms of the difference in $E_{3}$ estimates, $\sigma_{0}$, would have been $\sqrt{2}\,\sigma_{0}$ of the rms of individual $E_{3}$ estimates from 0.5 hr segments, i.e. 2 $\sqrt{2}\>\cdot\>1.42=2.01$ nrad. But we got $\sigma=1.47$ nrad, which is substantially less. We can compute correlation $\rho$ between estimates of $E_{3}$ angles derived from 0.5 hr date segments using a simple expression
$\rho=1-\frac{\displaystyle 1}{\displaystyle 2}\frac{\displaystyle\sigma^{2}}{\displaystyle\sigma_{0}^{2}}.$ (11)
This gives us $\rho=0.46$. We consider this correlation as a direct evidence of presence of a non-thermal noise due to mismodeling atmospheric path delay. Atmospheric path delay between two 0.5 hr time intervals is not independent and is correlated. Correlation in the atmospheric path delay between two segments causes correlation in estimates of $E_{3}$ angles determined from data of these segments. We should note that the correlation in the atmospheric path delay does not have to be the same as the correlation in $E_{3}$ estimates.
6.4 Formal and informal errors: what did we learn?
Our study confirmed what was already known for two centuries: uncertainties from ground astronomical observations computed using the error propagation law are wrong. But why? Gauss-Markov theorem states the condition when uncertainties derived by the formula of the error propagation law (that is why they are called formal) are the dispersion of estimated parameters. We have no doubts in the proof of Gauss-Markov theorem. Therefore, we interpret a digression of our accuracy estimates from formal errors as evidence of a violation of conditions of the Gauss-Markov theorem. Namely, we conclude that the weight matrix we used in data analysis differs from the invert of the covariance matrix of noise.
It was long known that atmospheric turbulence cannot be described as an uncorrelated random process [see, for in instance, 36]. However, a quantitative assessment of the impact of an incorrect stochastic model used in data analysis requires significant resources for acquiring redundant data and their analysis.
We see that formal errors derived from data analysis using weight matrices with zero off-diagonal terms, which is equivalent to an assumption of no correlation, cannot be used as a measure of errors in EOP estimations and cannot be used as a basis for optimization. Judging on formal errors, one can conclude that construction of new antennas at Wettzell and Kokee Park, installing next generation quad-band recording system improved the accuracy of $E_{3}$ angle by a factor of 2.1. Based on this metric one could make certain recommendations for the future development aimed at improving the EOP accuracy. However, considering the achieved change in accuracy, 15%, one could make rather different recommendations.
Our analysis demonstrated that at networks of size of 0.6–0.8 Earth diameter, atmospheric conditions are the primary factors that affect accuracy, other factors as bandwidth, number of observations per hour being secondary. We interpret it as a manifestation of the dominance of atmospheric noise. Our tests with an injection of the random noise demonstrate quantitatively that it is sufficient to reach a formal uncertainty of 20 ps of group delay determination for geodetic applications of VLBI techniques. Further improvement in group delay uncertainty does not buy us anything. For instance, EOP accuracy from quad-band vo observations with 4 ps group delay precision turned out even worse than EOP accuracy from dual-band r1r4 observations with 20 ps group delay precision. That does not necessarily mean that a change in VLBI hardware caused degradation in EOP accuracy. We should note the vo network was evolving in 2019–2024, and the scarcity of meridional baselines in the first part of the interval was the factor that affected EOP estimates.
As we see, the use of the additive weight update did not solve the problem of underestimation of $E_{3}$ angles following the error propagation law, but only alleviated the problem. Luati and Proietti [12] proved a theorem stating that least squares with a full a priori covariance matrix is equivalent to least squares with a diagonal weight matrix if and only if certain conditions are met, and weights and observation equations that we use in our solutions do not satisfy these conditions.
6.5 Observations of different durations
If the dependence of EOP accuracy on session duration were described by the power law -0.5, the distribution of long and short VLBI observing blocks over a given interval would not have mattered The broken power law of the dependence of EOP accuracy implies that planning observing sessions longer than 2–4 hours we get a diminishing return on used resources. Extending observing sessions from 4 to 24 hours, we have to spend 6 times more resources, but we gain only 70% in accuracy instead of 140%, because the atmospheric turbulence added a correlated noise. Adding more stations does not improve noticeably the accuracy of $E_{3}$ determination.
If $E_{3}$ were the only valuable data product of geodetic VLBI, then we recommend to limit observations to one long baseline of 0.6–0.8 Earth diameter but restrict observing session durations to several hours, and to run them several times a day. Considering other data products, the recommendation would change, however, collected evidence shows that running experiments for 24 hours for EOP time series estimation is far from optimal.
6.6 Comparison with IERS C04 time series
It is customary to compare EOP estimates against IERS C04 series [2]. Therefore, it is instructive to compare our estimates of EOP accuracy from analysis of concurrent VLBI experiments against the C04 time series. The time series of $E_{1}$ and $E_{2}$ angles in C04 time series are dominated by the IGS final data product from processing independent GNSS phase and pseudorange data. Therefore, one can expect that the statistics of the rms of differences against C04 will be close to our estimates of accuracy based on the three corner method. Indeed, it turned out very close for vo campaigns, but 13–17% higher for r1r4 campaigns. Time series $E_{3}$ in C04 are derived from the results of processing the same VLBI group delays. According to Bizouard et al. [2], the time series of $E_{3}$ angles are derived from a) $E_{3}$ estimates from analysis of R1 and R4 VLBI experiments by 10 analysis centers and then averaged with some weights by the so-called International VLBI Service for Geodesy and Astrometry (IVS) combination center; b) $E_{3}$ estimates from some 1-hr programs: 4×4$\times$4 × daily from the Institute of Applied Astronomy in Sankt-Peterburg, Russia, and several programs processed by the United States Naval Observatory (USNO) on days when there is no r1r4 experiment. The input time series are then smoothed by using the low-pass filter. Therefore, the IERS C04 accumulates errors due to deficiencies in data analysis of individual groups and suppresses, to some degree, fast variability in EOPs due to smoothing.
Both the differences of EOP estimates against IERS C04 and statistics of the difference from concurrent VLBI experiments provide measures of EOP accuracy. It is instructive to compare them. Table 9 shows these accuracy estimates.
Table 9: The estimates of EOP accuracy from the comparison of concurrent VLBI observations (column 3) and the rms of differences of EOP derived from our VLBI solution against IERS C04 (column 4). The estimates in column 3 for the first four rows are derived from the three-corner hat method, and for the last for rows are derived from comparison of EOP from concurrent observations.
Comp duration concur IERS C04
hour nrad nrad
$E_{1}$ vo 24 0.46 0.52
$E_{2}$ vo 24 0.57 0.67
$E_{1}$ r1r4 24 0.34 0.34
$E_{2}$ r1r4 24 0.32 0.35
$E_{3}$ vo/r1r4 24 0.44 0.74
$E_{3}$ K2Ws 1 1.30 1.52
$E_{3}$ MkWz 1 1.21 1.60
$E_{3}$ KkWz 1 1.50 1.48
It turned out that the differences of $E_{3}$ angles from 24-hr experiments are a factor of 1.7 greater than the accuracy of these angle estimates derived from our processing concurrent vo/r1r4 experiments (compare columns 3 and 4 in Table 9). In general, the accuracy of a combined series like IERS C04 is governed by the accuracy of input series and smoothing. We can rule out smoothing errors since the epochs of our EOP estimates and the EOP estimates of IVS combined series are practically the same. Therefore, we conclude that the IVS combined $E_{3}$ angle series are a factor of 1.7 less precise than our results. The differences of $E_{3}$ angles derived from 1-hr experiments with respect to the C04 series are a factor of 1.17–1.32 greater than the accuracy of these angle estimates from concurrent experiments, except for the time series from KkWz baselines.
In order to learn more, we took the IVS combined EOP series [^4] https://www.ccivs.bkg.bund.de/data/QUAT/COMBI/ivs24q3e.eops accessed on May 16, 2025.. We computed the statistics of the differences of IVS Combined EOP from 148 r1r4 experiments with respect to our solution after removal of the linear trend: the rms of the differences were 0.31, 0.38, and 0.86 nrad for $E_{1}$, $E_{2}$, and $E_{3}$ respectively. Similar statistics for EOP derived from 148 vo experiments are 0.61, 0.66, and 0.99 nrad.
To dig even deeper, we compared EOP estimates from the same list of concurrent vo and r1r4 observing programs using the IVS combined time series. The epochs of EOP in the IVS combined time series are close, within 2 hours, but not the same. We propagated the EOP on the common epoch using provided estimates of EOP rates. The rms of the differences divided by $\sqrt{2}$ are shown in Table 10. We see that our solution provides the rms of the differences a factor of 1.3 smaller for the $E_{3}$ angle. That prompts us to interpret the relatively large differences in IVS Combined EOP with respect to our solution as a deficiency of the IVS Combined series.
Table 10: The estimates of EOP accuracy from comparison of concurrent VLBI vo and r1r4 observing campaigns: our solution (the 2nd columns) and the IVS combined EOP time series (the 3rd column). The rms of the differences is divided by $\sqrt{2}$.
Comp Our solution IVS Combined
nrad nrad
$E_{1}$ 0.52 0.68
$E_{2}$ 0.71 0.69
$E_{3}$ 0.44 0.57
6.7 IVS Combined EOP time series and IERS C04
We established earlier that the IVS Combined EOP series has a notable deficiency. Since the time series from r1r4 experiments are used for IERS C04, it inherits this deficiency. We should note that the IVS combined solutions fall into the category of operational solutions, while our solutions fall into the category of re-processing. We base our analysis on 50 years of efforts to improve the VLBI data analysis technique pursued by NASA Goddard Space flight center, and our data analysis technique differs from the technique commonly used by other groups for operational data analysis. In particular, we see no biases, no systematic patterns when we compare EOP from 24-hr and 1-hr experiments or EOP from different observing sessions.
Computing a combined solution, i.e. the weighted mean of results, may have some advantage for operational time series because it helps to detect a failure in quality control of an individual analysis center and allows us to mitigate situations when a given analysis center is offline. On the other hand, processing the same dataset several times and then averaging results contradicts Gauss-Markov theorem, and that causes a deviation of such a result from the optimum.
We should bear in mind worse accuracy of a combined data series, treat it as a temporary, coarse solution, and avoid using it for science. IERS C04 time series falls into the same category of operational data products since it is based on operational solutions. We have demonstrated that EOP estimates from a uniform reprocessing using advanced theoretical model and parameter estimation technique of all accumulated data are superior. In a similar way, campaigns of reprocessing GNSS observations with an improved model [see, for instance, 3] also provide superior results.
6.8 Similar comparisons made by others
An idea of evaluation of $E_{3}$ accuracy from 24-hr and 1-hr VLBI experiments is not new. A commonly used approach is to compare $E_{3}$ against IERS C04 or similar combined series. For instance, Haas et al. [9] compared $E_{3}$ derived from twelve 1-hr experiments spanned over three months against the USNO Bulletin B. They presented rms of the differences from 11 solutions in a 1.7–2.3 nrad range (their Table 2). Schartner et al. [34] made a comparison of $E_{3}$ angle estimates from several 1 hr Intensive experiments are reported rms differences against JPL EOP2 and IERS C04 time series (their Tables 1 and 2) in a range of 1.3–5.6 nrad. Wang and Xu [42] computed the rms of differences in $E_{3}$ angles determined from 1-hr observing sessions at KkWz baseline and IERS C04 experiments over 2000–2011: 1.86 nrad. Yao et al. [44] compared $E_{3}$ derived from 61 1-hr observing sessions at baselines KkWz and K2Ws in 2021. They made comparison of $E_{3}$ angles with respect to IERS C04 and reported rms of the differences: 2.05 and 1.83 nrad. In addition, they split 1-hr K2Ws into two 0.5 hr long blocks in the beginning and the end of an 1-hr observing session and processed them individually. Yao et al. [44] report that the rms of differences with respect to IERS C04 increased by 7% when 0.5-hr blocks at the beginning of the observing session were processed and decreased by 10% when 30 minute blocks at the end of the observing sessions were processed. According to the F-test, these differences are insignificant at the 0.1 confidence level. That means that throwing away 1/2 of K2Ws data did not change results. This finding contradicts to the broken power law that we derived by thinning 24-hr datasets. Inspired by these results, we performed our analysis. Our analysis showed that throwing away half of the data degraded accuracy, but only at 9% instead of 41% if the noise were uncorrelated.
Raut et al. [23] compared $E_{3}$ angle estimates determined from processing 1-hr and 24-hr experiments at three independent campaigns during 15 days in November/December 2017. They performed a number of solutions, including splitting 24-hr experiments into 1-hr blocks that they called pseudo-intensives. They report in Table 5 the rms of differences with respect to IERS C04 1.97 nrad for 24-hr blocks and 1.85 nrad for 1-hr blocks. This finding also does not support the broken power law we found, but the rms of the differences are substantially greater than in our results.
For completeness, we mention that Gipson and Baver [7] published results of a campaign of 9 experiments with a design similar to our campaign s22. According to the campaign description, each experiment contained twenty two 1-hr scheduling blocks at KkWz baseline among with data at other stations. However, database files in the public archive at NASA Crustal Dynamics Data Information System (CDDIS) contains data at all stations, except stations kokee and wettzell. Apparently, Gipson and Baver [7] based their analysis on additional datasets that were not publicly released. Considering governing open science policies, we defer comments on these results till missing datasets will become available.
All these comparisons are made with respect to $E_{3}$ time series derived from other VLBI results with applying smoothing, which makes their interpretation difficult since three factors affect the differences, 1) intrinsic errors of $E_{3}$ from 1-hr experiments; 2) intrinsic errors of errors of $E_{3}$ that were used for generation of derived time series; and 3) errors introduced by smoothing.
MacMillan [13] presented wrms of the differences of EOP from two concurrent VLBI networks during 15 consecutive days during CONT17 observing campaign in November/December 2017. He reports the rms of the differences in Euler angle vector (0.21, 0.19, 0.25) nrad. This is a factor of 2–3 smaller than the rms of the differences in EOP derived from analysis of group delays from vo/r1r4 experiments presented in Table 4. There are three factors that played the role. First, CONT17 experiments are not representative. This campaign was an attempt to get the best results using existing VLBI hardware, while vo and r1r4 experiments have been running as regular, operational observing sessions. Second, these experiments ran in winter time for most of the stations in the network. And third, statistics over a continuous 15 day period are different from statistics over a 5 year period in a presence of red noise. Ray et al. [25] argues strongly that the longer the time period for error characterization, the greater the errors, and we find their argumentation quite convincing. Petrov and Kovalev [22] presented strong evidence of the impact of red noise on results of comparisons of two subsets of VLBI data on source positions. At present, we are not in a position to deliver a quantitative proof of the presence of red noise in EOP estimates at scales of 0.01–5 years, but we consider that the first two factors are not sufficient to explain the discrepancies in statistics of that magnitude.
7 Summary and Outlook
Having processed a number of recent and historical datasets of concurrent VLBI observations at different networks, we have arrived at more reliable estimates of the accuracy of time variable Euler angles describing the Earth’s rotation. These estimates have a little resemblance to formal computed using the law of error propagation under the assumption that the noise is uncorrelated. We summarize the accuracy of determination of $E_{3}$ angles from various 24-hr and 1-hr campaigns in Figure 11.
Figure 11: Accuracy of $E_{3}$ angle estimates derived from analysis of different concurrent experiments. The blue thick lines (leftmost 9) show accuracy from 1-hr experiments or processing block. The green thick lines (rightmost 2) show accuracy from 24-hr experiments.
We confirmed that the accuracy of EOP estimates from VLBI observations cannot be characterized by formal errors. Accuracy is determined mainly by the network and atmospheric processes.
We found strong evidence that the dominant error source that determines accuracy of geodetic VLBI results is the residual atmospheric path delay, which is variable and correlated. The dominance of the atmospheric noise manifests itself in a form of a) seasonality in EOP accuracy — winter results are accurate at the high confidence level; b) dependence of accuracy on experiment duration as the power law -0.3; c) the insensitivity of results on the precision of group delay when it drops below the 20 ps level; and d) correlation 0.46 between estimates in $E_{3}$ angles computed from independent group delay datasets 0.5 hours apart.
We found no evidence that advanced scheduling strategies for improved determination of the atmospheric path delay in zenith direction have any impact on the accuracy of EOP estimates. Inability to predict the accuracy based on simulation suggests that our approaches to data analysis and simulation of geodetic VLBI have to be revised. We recognize that this inability poses a serious problem. The use of a weight matrix with filled diagonal terms and zero off-diagonal terms was a necessity in the 19th and 20th centuries due to the scarcity of computing resources, but the use of this technique by inertia in the 21st century is an anachronism. We are going to enhance our data analysis procedure by the use of a full weight matrix, including off-diagonal terms, and encourage others to pursue these efforts. The goal of these efforts is to predict accuracy of EOP estimates within tens of percent, explain the power law -0.3 and seasonality in EOP accuracy, and explain correlation 0.46 between estimates of $E_{3}$ angles from 0.5-hr data segments. The dataset presented in this study is an excellent testbed that can be used for tuning algorithms that eventually are supposed to produce correct estimates of uncertainties of estimated parameters in the presence of a correlated noise.
The dependence of EOP errors as a power law of -0.3 for VLBI experiment durations longer than 4 hours and insensitivity of estimates of $E_{3}$ angle to the number of observing stations raises the issue of optimization of the program of geodetic VLBI observations. The current strategy of allocation of very significant resources for observing 24-hr sessions at a 10–12 station network with long gaps between sessions does not favor $E_{3}$ angle determination. The optimization of the observing strategy based on the power law of the dependence of EOP accuracy on session duration is a topic for a future research.
Although we have arrived at our conclusions based on the analysis of VLBI observations, the same atmospheric noise affects GNSS observations and therefore, our findings can be generalized to other microwave observations.
[…]
Acknowledgements
We would like to thank Frank Lemoine for valuable comments. It is our pleasure to thank staff at the CDDIS for making available Level 2 VLBI data. We thank Cynthia Thomas and Dirk Behrend for coordinating International VLBI for Geodesy and Astrometry observing programs.
Declarations
  • Funding: L.P. was supported by the NASA Space Geodesy Project.
  • Conflict of interest/Competing interests: none
  • Code availability: we used for our analysis open source software package SGDASS available at https://doi.org/10.25966/7z5f-eg07. The s22 schedules are generated using the VieSched++ software package [31] available at https://github.com/TUW-VieVS/VieSchedpp under the GNU General Public License.
  • Materials availability: all Level 2 VLBI data used in our analysis (group delays) are publicly available at https://cddis.nasa.gov/archive/vlbi/ivsdata
  • Author contribution: L.P., M.S., and C.P. designed s22 observing campaign; M.S. developed s22 experiment schedules; C.P. ran experiments, processed Level 0 and Level 1 data; L.P. ran data analysis of Level 2 data and performed statistical analysis of results. All authors discussed the results, commented on the manuscript, and agreed to be held accountable for the work.
  • Use of artificial intelligence: we certify that no machine learning or artificial intelligence techniques were used, neither during data analysis, nor in manuscript preparation.